OpenStudy (anonymous):

4. What is the greatest common factor of 42, 126, and 210 ? F. 2 G. 6 H. 14 J. 21 K. 42

OpenStudy (freckles):

I would write prime factorization for each. For example, 210=3(70)=3(7)(10)=3(7)(2)(5) With factors in order 210=2(3)(5)(7)

OpenStudy (freckles):

write the prime factorizations for each number

OpenStudy (freckles):

so we can compare what factors they have in common

OpenStudy (anonymous):

That's what I'm confused about. I'm very rusty on this stuff

OpenStudy (cwrw238):

divide by the lowest prime number then next prime number etc i'll give you an example prime factors of 18:-